last summer gary trained 32 more dogs than zina. together they trained 126 dogs. How many dogs did gary train?
Answer by stanbon(75887)

last summer gary trained 32 more dogs than zina. together they trained 126 dogs. How many dogs did gary train?
-----
g - z = 32
g + z = 126
-------------------
Add and solve for "g":
2g = 158
g = 79 (# trained by gary)
z = 126 - 79 = 47 (# trained by zina)
--------------
Cheers,
Stan H.
--------------